Scores for a common standardized college aptitude test are normally distributed with a mean of 484 and a standard deviation of 113. Randomly selected men are given a Test Preparation Course before taking this test. Assume, for sake of argument, that the preparation course has no effect. If 1 of the men is randomly selected, find the probability that his score is at least 582.2. A)P(X > 582.2) =

(Enter your answer as a number accurate to 4 decimal places.)

If 7 of the men are randomly selected, find the probability that their mean score is at least 582.2.

B)P(M > 582.2) =

(Enter your answer as a number accurate to 4 decimal places.) Assume that any probability less than 5% is sufficient evidence to conclude that the preparation course does help men do better. If the random sample of 7 men does result in a mean score of 582.2, is there strong evidence to support the claim that the course is actually effective? choose a or b

(a)Yes. The probability indicates that it is (highly ?) unlikely that by chance, a randomly selected group of students would get a mean as high as 582.2.

(b)No. The probability indicates that it is too possible by chance alone to randomly select a group of students with a mean as high as 582.2.
Answer #1

a)

µ =    484          
σ =    113          
              
P ( X ≥   582.2   ) = P( (X-µ)/σ ≥ (582.2-484) / 113)      
= P(Z ≥   0.87   ) = P(Z<-0.87 ) =    0.1924   (answer)

b)

n=   7          
              
X =   582.2          
Z =(X - µ )/(σ/√n) =    (582.2-484)/(113/√7)=       2.30  
P(X>582.2) = P(Z >2.3) = P(Z < -2.3 ) =           0.0107   (answer)

c)

(a)Yes. The probability indicates that it is (highly ?) unlikely that by chance, a randomly selected group of students would get a mean as high as 582.2.